# Regularity Priors for the Linear Atomic Cluster Expansion

### v1

@@ Abstract
Machine-learned interatomic potentials enable large systems to be simulated for long time scales at near ab-initio accuracy. This accuracy is achieved by fitting extremely flexible model architectures to high quality reference data. In practice, this flexibility can cause unwanted behavior such as jagged predicted potential energy surfaces and generally poor out-of-distribution behavior. We investigate a general strategy for incorporating prior beliefs on the regularity of the target energy into linear ACE models and explore to what extent this approach improves the quality of the fitted models. Our main focus is an over-regularisation that replicates the Gaussian broadening used in SOAP descriptors within the ACE framework. Numerical tests indicate that the exact form of the prior is non-critical but that including such a prior leads to significant improvement in test errors, consistent repulsion at close-approach, eliminates spurious false minima in the potential energy and enhances stability during molecular dynamics simulations.
@@ I Introduction
However, the extreme flexibility of ML architectures has drawbacks. In particular, MLIPs are notoriously poor at extrapolating outside the domain of the training dataset [vita2023data, benoit2020measuring], which often leads to catastrophic failure during Molecular Dynamics simulations e.g. a “hole” in the Potential Energy Surface is encountered, the predicted energy decreases by $\sim$100s-1000s of eV/atom and the simulated system explodes [nandi2019using, van2020regularised]. Avoiding such behavior tends to require careful construction of the training dataset, either using labour intensive manually chosen configurations, or via automated approaches such as active learning [van2023hyperactive, sivaraman2020machine]. The recent wave of foundation potentials [batatia2023foundation, yang2024mattersim, neumann2024orb, park2024scalable, fu2025learning, barroso2024open, merchant2023scaling, rhodes2025orb, bochkarev2024graph], have come a long way towards fixing this issue, primarily by using enormous training datasets in combination with careful architectural decisions. These potentials show unprecedented stability across a diverse range of materials and simulations [batatia2023foundation] whilst retaining meaningful accuracy. However, they are not perfect, with the occasional hole still encountered for particularly unusual configurations or, more commonly, for high pressure configurations under extreme compression where there is relatively little training data available. Furthermore, some architectures are capable of achieving excellent point wise errors on energies, forces and stresses [riebesell2025framework] whilst interpolating between the training data in a surprisingly “rough” fashion, with the predicted Potential Energy Surface showing small amplitude high frequency oscillations, which can be sufficiently sharp to appear almost as kinks. This behavior is seen in the homo-nuclear diatomic energy curves presented in the MLIP Arena benchmark [chiang2025mlip] and can lead to many practical concerns, such as failure of geometry optimization or transition path search [henkelman2000climbing] or encountering false minima during crystal structure prediction [deringer2018data].
@@
Figure 1: The energy and force predictions on the Mg-Mg dimer are shown for various foundation models [batatia2023foundation, yang2024mattersim, neumann2024orb, park2024scalable, fu2025learning, barroso2024open] as a function of dimer separation. The MLIP Arena python package [chiang2025mlip] was used to load all models and obtain the reference PBE result, originally computed with VASP [kresse1996efficiency]. For the avoidance of doubt model versions are MACE-MPA-0, MatterSim-v1.0.0-5M, ORB-v2, 7net-0, eSEN-30M-OAM and eqV2_86M_omat_mp_salex.
@@
The issues described above are “obvious” failings in the sense that we can identify them almost at a glance. This is because i) empirically, the reference Potential Energy Surface in most low-energy regions of configuration space for chemical and materials systems is smooth in the sense that it varies over a characteristic length scale that depends on the size of the atoms in questions, for example see Figure 1, ii) the energy should increase sharply at short interatomic distances $r$ due to a combination of Pauli and Coulomb repulsion, with the $1/r$ Coulomb repulsion between nuclei dominating at very close approach [ziegler1985stopping, nordlund2025repulsive]. In this work we investigate how this chemical intuition can be incorporated into linear ACE models by adjusting the form of the regularization term. We focus on ACE for two reasons: (i) it remains an attractive light-weight alternative [zhou2025full] to large-scale message passing neural networks; (ii) it provides a prototype study that can be extended to many related architectures, such as MACE [batatia2022mace] and GRACE [bochkarev2024graph], that build on the ACE framework.
@@
When fitting linear models the standard choice is to employ Tikhonov regularization. This leads to a single parameter, the regularization strength, which balances model accuracy (on the training set) against stability, by penalizing models with large coefficients. From a Bayesian perspective, such a choice implies a uniform prior belief for the coefficients of all basis functions, irrespective of their polynomial degree. Here, we investigate different “shapes” for the regularization which aim to incorporate our prior belief on the smoothness of the PES. By smoothness, or more formally regularity, we not only mean that the energy should be continuous and differentiable, but also that it should not contain small amplitude high-frequency oscillations. This is analogous to the length-scale parameter within Gaussian Process Regression kernels [bartok2010gaussian, deringer2021gaussian] and we show how the $\sigma$ parameter in SOAP descriptors [bartok2013representing] corresponds to a particular choice of regularization within linear ACE. In practice, we find the exact form of the regularization term is not critical, but that picking one which incorporates a notion of regularity leads to significant improvement in model performance without incurring any additional computation cost. We validate this with a variety of numerical tests including error estimates, Molecular Dynamics stability, Random Structure Searching and comparing various cuts through the predicted Potential Energy Surface with the true reference.
@@ II.1 Regularized Atomic Cluster Expansion
We are foremost concerned with different choices for $\Gamma$, how they can be interpreted and how they affect model performance and predicted properties. Specifically, we will explain how certain choices of $\gamma$ can be interpreted as regularity priors for the potentials $U_{N}$ and, by investigating the effect of each prior on the implied atomic neighbor density, draw a connection to the Gaussian broadening used in SOAP-GAP potentials.
@@ II.2 Regularity Priors
The regularity class of functions having the analogous decay of coefficients is extremely small (the algebra of polynomials and Gaussians), hence this choice should be thought of as over-regularization and not as describing the regularity of the underlying target function. A detailed dicussion on the interplay between regularity priors and symmetrising over the rotation group is given in section VII.1.
@@ II.4 Implied Neighbor Density
ACE features can be understood in terms of symmetrized tensor products of a fictitious atomic neighbor density formed by representing each neighboring atom as delta function. The neighbour density can be written as
@@
where here the $A_{i,nlm}$ are expansion coefficients. As such, scaling the expansion coefficients according to a specific regularity prior $\tilde{A}_{i,n_{t}l_{t}m_{t}}=\gamma_{n_{t}l_{t}}^{-1}A_{in_{t}l_{t}m_{t}}$ can be interpreted as modifying the corresponding neighbour density, $\tilde{\rho}$. This connection is particularly interesting as GAP potentials, which use $N=2$ ACE features where each atom is represented by a Gaussian of width $\sigma$, have empirically shown better stability than ACE potentials fit to the same dataset[morrow2022indirect]. To examine this relationship further, let us assume that $\phi_{nlm}$ are Laplacian eigenfunctions, i.e., are $L^{2}$-orthonormal, satisfy $\Delta\phi_{nlm}=\lambda_{nl}\phi_{nlm}$ and $\phi_{nlm}(|{\bm{r}}|=r_{\mathrm{cut}})=0$. Note, $\lambda_{nl}$ has no $m$ dependence as $\Delta$ is isotropic. Upon writing $\gamma_{nl}^{-1}=f(\lambda_{nl})$, we then expand a general modified density
@@
In principle, any operator could be used in place of $\Delta$ but, as noted by Bigi et. al. [bigi2022smooth], the Laplacian is a natural choice as, subject to orthonormality constraints, its eigenstates are constructed to minimize their average squared gradient, with $\lambda_{nl}$ quantifying this measure of smoothness. Importantly for our context, powers of $\Delta$ provide the means of specifying regularity classes. For example, one could minimize curvature by considering eigenfunctions of the Bi-Laplacian $\Delta^{2}$ or one could measure smoothness with respect to a scaled radial coordinate. Here, we choose to use $\Delta$ as i) $\lambda_{nl}$ scales approximately as $\sim n^{2}+l^{2}$, which allows for a simple correspondence with our regularity priors and ii) it enables approximate expressions for $\tilde{\rho}$ to be derived. This allows us to better understand the effects of the different priors and make a connection to the Gaussian neighbour density used in SOAP.
@@
which implies the associated neighbor density $\tilde{\rho}=H_{\alpha}\ast\rho_{i}^{(0)}$ has a Lorentzian shape. A comparison between the smoothing effects of all three priors on $\tilde{\rho}$ is shown in Figure 2 for $n_{\mathrm{max}}=l_{\mathrm{max}}=20$. Interestingly, despite the differences between the forms of the priors, the actual neighbor densities are strikingly similar once a moderate level of smoothing is applied.
@@
III Results
The ACEPotentials.jl package was used to fit all ACE models shown in the subsequent section. This package uses a radial basis of the form
@@ III Results
where $y_{ij}=y(Z_{i},Z_{j},r_{ij})$ denotes a coordinate transform, $f_{\rm env}$ is an envelope function enforcing predefined behavior as $r\to 0$ and $r\to r_{\rm cut}$, $P_{n}$ are shifted and scaled Legendre polynomials, and the chemical element dependence is folded into the composite index $n=(n^{\prime},\zeta)$. Explicit functional forms are provided in [witt2023acepotentials]. The regularity priors are therefore to be understood as acting in the transformed $y$-space and after factoring out the envelope function. The priors were applied using the definitions in (8) and (10) with $\sigma_{\mathrm{l}}=1.5\sigma_{\mathrm{n}}$ used for the both the exponential and algebraic priors, which corresponds to the default $w_{L}$ parameter used to determine basis set truncation [witt2023acepotentials]. For the Gaussian prior $\sigma_{\mathrm{n}}=\sqrt{5}\sigma_{\mathrm{l}}$, was used, with this ratio determined by the expansion of a Gaussian in the Laplcaian Eigenfunctions basis of ref. [bigi2022smooth], see section VII.2 for details.
@@ III.1 Silicon
The Silicon-GAP-18 dataset from [bartok2018machine] was used to investigate the practical consequences of fitting ACE potentials using various regularity priors. This database is very structurally diverse, containing a mixture of crystalline, amorphous and liquid configurations as well as various crystalline surfaces, defects and dislocations. This diversity ensures good coverage of the silicon potential energy surface and enabled the authors to fit a general-purpose potential [bartok2018machine] using the Gaussian Approximation Potential [bartok2010gaussian, deringer2021gaussian, klawohn2023gaussian] framework. Here we fit ACE potentials to the original dataset and a smaller subset, referred to here as Si10pc, containing $\sim$10% of the original data. We focus on results obtained on Si10pc, where the differences between different choices of priors should be more pronounced. Recent work has shown that a carefully selected 5% of the original dataset is sufficient to fit high-quality ACE potentials [van2023hyperactive]; by contrast, Si10pc was constructed by applying a random 10:90 train:test split to the full dataset where the split was stratified by both configuration size and type. Additionally, the most and least dense configuration of each type was included in the training set to avoid excessive extrapolation in density.
@@
Figure 3: Force and energy errors on configurations in the training and test sets, as well as two different configuration types from ref. [morrow2022indirect], see main text for details, not present in the training data, are shown for ACE models fit using Gaussian regularity priors. Larger values of sigma correspond to a more regular prior. The force error on the training set, shown in the legend, is constant as this was used to select the strength of the regularization.
The ACE potentials were constructed using a local cutoff of $r_{\rm cut}=5$$\text{\AA }$, correlation order $N_{\rm max}=4$ (5-body) features and a maximum polynomial degree of 21 for all correlation orders.
To promote repulsion at very short inter-atomic distances, where Coulomb repulsion between nuclei dominates [ziegler1985stopping, nordlund2025repulsive], the training data was augmented with a single synthetic dimer configuration labeled with energy $E_{\mathrm{syn}}$=100 eV and with separation $r_{\mathrm{syn}}$, defined by $f_{\mathrm{env}}(r_{\mathrm{syn}})=E_{\mathrm{syn}}$ for each pair of elements. Observations of energies, forces and stresses were given relative weights of $w_{E}=30$, $w_{F}=1$ and $w_{S}=1$ respectively with much smaller weights of $w_{\mathrm{syn}}=0.01$ used for the synthetic dimer data.
@@
Figure 4: Force RMSE on the training and test sets is shown for potential fit to Si10pc using a Gaussian regularity prior with varying $\sigma$. Each marker corresponds to a particular choice of Tikhonov regularization strength.
@@
The potentials were then fit using simple Tikhonov regularization applied after the regularity prior was used to scale the basis functions appropriately as outlined in Section II.3. To distinguish between the strength and the “shape” of the regularization, i.e. different form of regularity prior, the regularization strength was tuned such that the force RMSE on the training set was equal for all models. This was repeated for three different target training force RMSEs with the resulting energy and force errors on the training and test sets shown in Figure 3 for models fit using the Gaussian regularity prior. The errors on two further configurations types, 100 configurations from Random Structure Search [pickard2011ab] and 100 configurations from a Molecular Dynamics simulation where amorphous Si was compressed from 0 to 20 GPa, were also evaluated. These configurations were taken from ref [morrow2022indirect] and are used to illustrate the effect of regularity priors in a highly extrapolative regime. For all configuration types, using $\sigma=2$ lead to $\sim$ 40% reduction in force RMSE and up to 80% reduction in energy RMSE compared to $\sigma=0$ (no regularity prior). As shown in Figure 4, this is not caused by the specific choice of target force RMSE, with these three values covering the minimum achievable test force RMSE for all values of $\sigma$. It is notable that despite the equivalence between the Gaussian regularity priors and SOAP features, outlined in Section II.3, the optimal value for $\sigma$ is much larger than would typically be used in SOAP-GAP potentials where $\sigma=0.3-0.5$ Å is more typical [deringer2018data]. This can be partially attributed to the distance transform applied to the radial coordinate, which increases the resolution near the typical nearest-neighbor position, effectively reducing $\sigma$ where there is most data and increasing it elsewhere. We also note that the exact shape of the regularity prior appears to make little difference to the achievable errors, with very similar performance achieved using both the algebraic and exponential regularity priors as shown in Figure S2. Given this, we present results for the Gaussian prior (Force RMSE on the training set = 0.075 eV/Å) only in the main text, with corresponding Figures for the algebraic and exponential priors provided in Figures S4 and S5 in the Appendix.
@@
Figure 5: The stress (derivative of energy) is plotted as a function of separation for decohesion of silicon diamond into the 100, 110 and 111 surfaces. The curves for different values of $\sigma$ are offset (black arrows) to aid comparison. The bottom right panel shows the energy of the Si-Si dimer as a function of separation between the atoms.
@@
Whilst reducing errors is highly desirable it is not the only important metric. For instance, many modern architectures achieve excellent errors but are unable to perform stable molecular dynamics simulations [fu2022forces] or, do so at the cost of having very rough potential energy surfaces with many spurious additional minima, see Figure 1. Quantifying this is challenging but we can probe it by simply plotting various cuts through the potential energy surface. The simplest example of this is the Si-Si dimer interaction energy which is shown in the bottom right of Figure 5. Despite all models having the same training force RMSE, the models with large values of $\sigma$ bear much closer resemblance to the reference Density-Functional Theory dimer curve. Crucially, they exhibit no false minima and are repulsive at close-approach, despite there being no dimer data in the training set.
@@
Inspecting higher body-order terms individually is both impractical and challenging to visualize due to the increased number of parameters. As such, we instead analyze them by constructing a series of 1D cuts through the full PES that depend on all body-orders present in the potential. The first example was taken from ref. [bartok2018machine] and involves computing the decohesion energy of the Si-diamond structure as it is rigidly separated to create a slab with two surfaces separated by vacuum. The derivative of the decohesion energy, which has the units of stress, is shown for trajectories that created (100), (110) and (111) surfaces in Figure 5. The reference DFT stress is notably smooth, with a single maximum and no further oscillations. In contrast, the $\sigma=0$ ACE potentials show clear spurious oscillations which are gradually damped away as $\sigma$ is increased. Smaller differences are observed in the decohesion energy itself, shown in Figure S1, with variations in $\sigma$ having little effect on the final surface energy; this is expected as these configurations are present in the training set. A similar test was performed whereby Si-diamond was rigidly exfoliated to individual layers of silicene. Here, clear differences in the energy, left panel of Figure 6, are seen between the potentials with all ACE models having an additional energy barrier to exfoliation that is not present in the Density-Functional Theory reference. The height of this barrier, and the size of associated oscillations in the derivative of the energy, decrease as $\sigma$ is increased.
@@
Figure 6: The energy (left) and its derivative with respect to added vacuum (right) are shown as the diamond structure silicon crystal is rigidly exfoliated into layers of silicene.
@@
Figure 7: The change in energy as an atom is move across the vacuum from the bottom (110) surface of a silicon slab to the top (110) surface is shown on the left. On the right, the z component of the force on the atom being moved is shown. Curves for different values of $\sigma$ are offset (black arrows) to aid comparison.

### v2

@@ Abstract
Machine-learned interatomic potentials enable large systems to be simulated for long time scales at near ab-initio accuracy. This accuracy is achieved by fitting extremely flexible model architectures to high quality reference data. In practice, this flexibility can cause unwanted behavior such as jagged predicted potential energy surfaces and generally poor out-of-distribution behavior. We investigate a general strategy for incorporating prior beliefs on the regularity of the target energy into linear Atomic Cluster Expansion (ACE) models and explore to what extent this approach improves the quality of the fitted models. Our main focus is an over-regularisation that replicates the Gaussian broadening used in Smooth Overlap of Atomic Positions (SOAP) descriptors within the ACE framework. Numerical tests indicate that the exact form of the prior is non-critical but that including such a prior leads to significant improvement in test errors, consistent repulsion at close-approach, eliminates spurious false minima in the potential energy and enhances stability during molecular dynamics simulations.
@@ I Introduction
However, the extreme flexibility of ML architectures has drawbacks. In particular, MLIPs are notoriously poor at extrapolating outside the domain of the training dataset [vita2023data, benoit2020measuring], which often leads to catastrophic failure during Molecular Dynamics simulations e.g. a “hole” in the Potential Energy Surface could be encountered causing the predicted energy to decrease by $\sim$100s-1000s of eV/atom and the simulated system to explode [nandi2019using, van2020regularised]. Avoiding such behavior tends to require careful construction of the training dataset, either using labour intensive manually chosen configurations, or via automated approaches such as active learning [van2023hyperactive, sivaraman2020machine]. The recent wave of foundation potentials [batatia2023foundation, yang2024mattersim, neumann2024orb, park2024scalable, fu2025learning, barroso2024open, merchant2023scaling, rhodes2025orb, bochkarev2024graph] have come a long way towards fixing this issue, primarily by using enormous training datasets in combination with careful architectural decisions. These potentials show unprecedented stability across a diverse range of materials and simulations [batatia2023foundation] whilst retaining meaningful accuracy. However, they are not perfect, with the occasional hole still encountered for particularly unusual configurations or, more commonly, for high pressure configurations under extreme compression where there is relatively little training data available. Furthermore, some architectures are capable of achieving excellent point wise errors on energies, forces and stresses [riebesell2025framework] whilst interpolating between the training data in a surprisingly “rough” fashion, with the predicted Potential Energy Surface showing small amplitude high frequency oscillations, which can be sufficiently sharp to appear almost as kinks. This behavior is seen in the homo-nuclear diatomic energy curves presented in the MLIP Arena benchmark [chiang2025mlip] and can lead to many practical concerns, such as failure of geometry optimization or transition path search [henkelman2000climbing] or encountering false minima during crystal structure prediction [deringer2018data].
Refer to caption(a)(b)
@@
Figure 1: The energy (a) and force (b) predictions on the Mg-Mg dimer are shown for various foundation models [batatia2023foundation, yang2024mattersim, neumann2024orb, park2024scalable, fu2025learning, barroso2024open] as a function of dimer separation. The MLIP Arena python package [chiang2025mlip] was used to load all models and obtain the reference PBE result, originally computed with VASP [kresse1996efficiency]. For the avoidance of doubt, model versions are MACE-MPA-0, MatterSim-v1.0.0-5M, ORB-v2, 7net-0, eSEN-30M-OAM and eqV2_86M_omat_mp_salex.
@@
The issues described above are “obvious” failings in the sense that we can identify them almost at a glance for two key reasons. Firstly, empirically, the reference Potential Energy Surface in most low-energy regions of configuration space for chemical and materials systems is smooth in the sense that it varies over a characteristic length scale that depends on the size of the atoms in questions, for example see figure1. Secondly,the energy should increase sharply at short interatomic distances $r$ due to a combination of Pauli and Coulomb repulsion, with the $1/r$ Coulomb repulsion between nuclei dominating at very close approach [ziegler1985stopping, nordlund2025repulsive]. In this work we investigate how this chemical intuition can be incorporated into linear Atomic Cluster Expansion models by adjusting the form of the regularization term. We focus on ACE for two reasons: (i) it remains an attractive light-weight alternative [zhou2025full] to large-scale message passing neural networks; and (ii) it provides a prototype study that can be extended to many related architectures, such as MACE [batatia2022mace] and GRACE [bochkarev2024graph], that build on the ACE framework.
@@
When fitting linear models the standard choice is to employ Tikhonov regularization. This leads to a single parameter, the regularization strength, which balances model accuracy (on the training set) against stability, by penalizing models with large coefficients. From a Bayesian perspective, such a choice implies a uniform prior belief for the coefficients of all basis functions, irrespective of their polynomial degree. Here, we investigate different “shapes” for the regularization which aim to incorporate our prior belief on the smoothness of the PES. By smoothness, or more formally regularity, we not only mean that the energy should be continuous and differentiable, but also that it should not contain small amplitude high-frequency oscillations. This is analogous to the length-scale parameter within Gaussian Process Regression kernels [bartok2010gaussian, deringer2021gaussian] and we show how the $\sigma$ parameter in Smooth Overlap of Atomic Positions descriptors [bartok2013representing] corresponds to a particular choice of regularization within linear ACE. In practice, we find that the exact form of the regularization term is not critical, but that picking one which incorporates a notion of regularity leads to significant improvement in model performance without incurring any additional computational cost. We validate this with a variety of numerical tests including error estimates, Molecular Dynamics stability, and Random Structure Searching as well as by comparing various cuts through the predicted Potential Energy Surface with the true reference.
@@ II.1 Regularized Atomic Cluster Expansion
We are foremost concerned with different choices for $\Gamma$, how they can be interpreted and how they affect model performance and predicted properties. Specifically, we will explain how certain choices of $\gamma$ can be interpreted as regularity priors for the potentials $U_{N}$ and, by investigating the effect of each prior on the implied atomic neighbor density, draw a connection to the Gaussian broadening used when combining Gaussian Approximation Potentials with SOAP descriptors[klawohn2023gaussian].
@@ II.2 Regularity Priors
The regularity class of functions having the analogous decay of coefficients is extremely small (the algebra of polynomials and Gaussians), hence this choice should be thought of as over-regularization and not as describing the regularity of the underlying target function. A detailed discussion on the interplay between regularity priors and symmetrization over the rotation group is given in section VII.1.
@@ II.4 Implied Neighbor Density
ACE features can be understood in terms of symmetrized tensor products of a fictitious atomic neighbor density formed by representing each neighboring atom as a delta function. The neighbour density can be written as
@@
where here the $A_{i,nlm}$ are expansion coefficients. As such, scaling the expansion coefficients according to a specific regularity prior $\tilde{A}_{i,n_{t}l_{t}m_{t}}=\gamma_{n_{t}l_{t}}^{-1}A_{in_{t}l_{t}m_{t}}$ can be interpreted as modifying the corresponding neighbour density, $\tilde{\rho}$. This connection is particularly interesting since GAP potentials, which use $N=2$ ACE features where each atom is represented by a Gaussian of width $\sigma$, have empirically shown better stability than ACE potentials fit to the same dataset [morrow2022indirect]. To examine this relationship further, let us assume that $\phi_{nlm}$ are Laplacian eigenfunctions, i.e., are $L^{2}$-orthonormal and satisfy $\Delta\phi_{nlm}=\lambda_{nl}\phi_{nlm}$ as well as $\phi_{nlm}(|{\bm{r}}|=r_{\mathrm{cut}})=0$. Note that $\lambda_{nl}$ has no $m$ dependence as $\Delta$ is isotropic. Upon writing $\gamma_{nl}^{-1}=f(\lambda_{nl})$, we then expand a general modified density
@@
In principle, any operator could be used in place of $\Delta$ but, as noted by Bigi et. al. [bigi2022smooth], the Laplacian is a natural choice as. Specificially, its eigenstates are constructed to minimize their average squared gradient, subject to orthonormality constraints, with $\lambda_{nl}$ quantifying this measure of smoothness. Importantly for our context, powers of $\Delta$ provide the means of specifying regularity classes. For example, one could minimize curvature by considering eigenfunctions of the Bi-Laplacian $\Delta^{2}$ or one could measure smoothness with respect to a scaled radial coordinate. Here, we choose to use $\Delta$ as i) $\lambda_{nl}$ scales approximately as $\sim n^{2}+l^{2}$, which allows for a simple correspondence with our regularity priors, and ii) it enables approximate expressions for $\tilde{\rho}$ to be derived. This allows us to better understand the effects of the different priors and make a connection to the Gaussian neighbour density used in SOAP.
@@
which implies the associated neighbor density $\tilde{\rho}=H_{\alpha}\ast\rho_{i}^{(0)}$ has a Lorentzian shape. A comparison between the smoothing effects of all three priors, i.e. polynomial, exponential and Gaussian, on $\tilde{\rho}$ is shown in figure 2 for $n_{\mathrm{max}}=l_{\mathrm{max}}=20$. Interestingly, despite the differences between the forms of the priors, the actual neighbor densities are strikingly similar once a moderate level of smoothing is applied.
Refer to caption(a)(b)(c)(d)(e)(f)
@@
II.5 Potential Fitting
The ACEPotentials.jl package [witt2023acepotentials] was used to fit all ACE models shown in the subsequent section. This package uses a radial basis of the form
@@ II.5 Potential Fitting
where $y_{ij}=y(Z_{i},Z_{j},r_{ij})$ denotes a coordinate transform, $f_{\rm env}$ is an envelope function enforcing predefined behavior as $r\to 0$ and $r\to r_{\rm cut}$, $P_{n}$ are shifted and scaled Legendre polynomials, while the chemical element dependence is folded into the composite index $n=(n^{\prime},\zeta)$. Explicit functional forms are provided in [witt2023acepotentials]. The regularity priors are therefore to be understood as acting in the transformed $y$-space and after factoring out the envelope function. The priors were applied using the definitions in (8) and (10) with $\sigma_{\mathrm{l}}=1.5\sigma_{\mathrm{n}}$ used for both the exponential and algebraic priors, since this corresponds to the default $w_{L}$ parameter for determining basis set truncation [witt2023acepotentials]. For the Gaussian prior a different ratio, $\sigma_{\mathrm{n}}=\sqrt{5}\sigma_{\mathrm{l}}$, was used; determined by the expansion of a Gaussian in the Laplcaian Eigenfunctions basis of ref. [bigi2022smooth], see section VII.2 for details. To better facilitate comparison to typical value used with Smooth Overlap of Atomic Positions descriptors, all values of $\sigma$ for the Gaussian prior will be quoted in Å where, $\sigma_{n}=\sqrt{5}\sigma/r_{\mathrm{cut}}$ and $\sigma_{l}=\sigma/r_{\mathrm{cut}}$.
To promote repulsion at very short inter-atomic distances, where Coulomb repulsion between nuclei dominates [ziegler1985stopping, nordlund2025repulsive], the training data was augmented with a single synthetic dimer configuration labeled with energy $E_{\mathrm{syn}}$=100 eV and with separation $r_{\mathrm{syn}}$, defined by $f_{\mathrm{env}}(r_{\mathrm{syn}})=E_{\mathrm{syn}}$ for each pair of elements. Observations of energies, forces and stresses were given relative weights of $w_{E}=30$, $w_{F}=1$ and $w_{S}=1$ respectively while much smaller weights of $w_{\mathrm{syn}}=0.01$ used for the synthetic dimer data. Furthermore, a purified version of the ACE basis was used, where the two-body contributions were removed from all higher order terms, see refs. [witt2023acepotentials, ho2024atomic] for a detailed discussion.
III Results
@@ III.1 Silicon
The Silicon-GAP-18 dataset from [bartok2018machine] was used to investigate the practical consequences of fitting ACE potentials using various regularity priors. This database is very structurally diverse, containing a mixture of crystalline, amorphous and liquid configurations as well as various crystalline surfaces, defects and dislocations. This diversity ensures good coverage of the silicon potential energy surface and made it possible o fit a general-purpose potential [bartok2018machine] using the Gaussian Approximation Potential [bartok2010gaussian, deringer2021gaussian, klawohn2023gaussian] framework. Specifically we fit ACE potentials to the original dataset and a smaller subset, referred to as Si10pc, containing $\sim$10% of the original data. We focus on results obtained on Si10pc, where the differences between different choices of priors should be more pronounced. Recent work has shown that a carefully selected 5% of the original dataset is sufficient to fit high-quality ACE potentials [van2023hyperactive]; by contrast, Si10pc was constructed by applying a random 10:90 train:test split to the full dataset where the split was stratified by both configuration size and type. Additionally, the most and least dense configuration of each type was included in the training set to avoid excessive extrapolation in density.
Refer to caption(a)(b)(c)(d)(e)(f)(g)(h)
@@
Figure 3: Force and energy errors are shown for ACE models fit using Gaussian regularity priors on configurations in the training and test sets, as well as two different configuration types from ref. [morrow2022indirect] that are not present in the training data, see main text for details . Larger values of $\sigma$ correspond to a more regularized prior. The force error on the training set, shown in the legend, is constant as this was used to select the strength of the regularization.
@@
Figure 4: Force RMSE on the training and test sets is shown for potential fit to Si10pc using a Gaussian regularity prior with varying $\sigma$. Each marker corresponds to a particular choice of Tikhonov regularization strength. The dashed black lines correspond to the three different training force RMSEs shown in figure 3.
@@
The ACE potentials were constructed using a $r_{\rm cut}=5$$\text{\AA }$ local cutoff and correlation orders up to $N_{\rm max}=4$ (5-body) features, with a maximum polynomial degree of 21 for all correlation orders. The potentials were then fit using simple Tikhonov regularization applied after the regularity prior was used to scale the basis functions appropriately as outlined in Section II.3. To distinguish between the strength and the “shape” of the regularization, i.e. different form of regularity prior, the regularization strength was tuned such that the force root mean square error (RMSE) on the training set was equal for all models. This was repeated for three different target training force RMSEs with the resulting energy and force errors on the training and test sets shown in figure 3 for models fit using the Gaussian regularity prior. Errors were also evaluated for two further configurations types: 100 configurations from Random Structure Searching[pickard2011ab] and 100 configurations from a Molecular Dynamics simulation where amorphous Si was compressed from 0 to 20 GPa. These configurations were taken from ref [morrow2022indirect] and are used to illustrate the effect of regularity priors in a highly extrapolative regime. For all configuration types, using $\sigma=2$ Å lead to $\sim$ 40% reduction in force RMSE and up to 80% reduction in energy RMSE compared to $\sigma=0$ (no regularity prior). As is evident from figure 4, which shows the error as a function of $\sigma$ for three different choices that cover the minimum achievable test force RMSE, this is achieved regardless of the specific target force RMSE It is notable that despite the equivalence between the Gaussian regularity priors and SOAP features, outlined in Section II.3, the optimal value for $\sigma$ is much larger than would typically be used in SOAP-GAP potentials where $\sigma=0.3-0.5$ Å is more typical [deringer2018data]. This can be partially attributed to the distance transform applied to the radial coordinate, which increases the resolution near the typical nearest-neighbor position, effectively reducing $\sigma$ where there is most data and increasing it elsewhere. We also note that the exact shape of the regularity prior appears to make little difference to the achievable errors, with very similar performance achieved using both the algebraic and exponential regularity priors as shown in figure S2. Given this, we only present results for the Gaussian prior (Force RMSE on the training set = 0.075 eV/Å) in the main text, with corresponding Figures for the algebraic and exponential priors provided in figures S4 and S5 in the Appendix.
Refer to caption(a)(b)(c)(d)
@@
Figure 5: (a)-(c))The stress (derivative of energy) is plotted as a function of separation for decohesion of silicon diamond into the (100), (110) and (111) surfaces. The curves for different values of $\sigma$ are offset (black arrows) to aid comparison. (d) The bottom right panel shows the energy of the Si-Si dimer as a function of separation between the atoms.
@@
Whilst reducing errors is highly desirable it is not the only important metric. For instance, many modern architectures achieve excellent errors but are unable to perform stable Molecular Dynamics simulations [fu2022forces], or do so at the cost of having very rough potential energy surfaces with many spurious additional minima, see figure 1. Quantifying this is challenging but we can probe it by simply plotting various cuts through the potential energy surface. The simplest example of this is the Si-Si dimer interaction energy which is shown in the bottom right of figure 5. Despite all models having the same training force RMSE, the models with large values of $\sigma$ bear much closer resemblance to the reference Density-Functional Theory dimer curve. Crucially, they exhibit no false minima and are repulsive at close-approach, despite there being no dimer data in the training set.
@@
Inspecting higher body-order terms individually is both impractical and challenging to visualize due to the increased number of parameters. As such, we instead analyze them by constructing a series of 1D cuts through the full PES that depend on all body-orders present in the potential. The first example was taken from ref. [bartok2018machine] and involves computing the decohesion energy of the Si-diamond structure as it is rigidly separated to create a slab with two surfaces separated by vacuum. The derivative of the decohesion energy, which has the units of stress, is shown in figure 5 for trajectories that created (100), (110) and (111) surfaces . The reference DFT stress is notably smooth, with a single maximum and no further oscillations. In contrast, the $\sigma=0$ ACE potentials show clear spurious oscillations which are gradually damped away as $\sigma$ is increased. Smaller differences are observed in the decohesion energy itself, shown in figure S1, with variations in $\sigma$ having little effect on the final surface energy; this is expected as these configurations are present in the training set. A similar test was performed whereby Si-diamond was rigidly exfoliated to individual layers of silicene. Here, clear differences in the energy are observed between the potentials with all ACE models having an additional energy barrier to exfoliation that is not present in the Density-Functional Theory reference; see the left panel of figure 6. The height of this barrier, and the size of associated oscillations in the derivative of the energy, decrease as $\sigma$ is increased.
Refer to caption(a)(b)
@@
Figure 6: The energy (a) and its derivative with respect to added vacuum (b) are shown as the diamond structure silicon crystal is rigidly exfoliated into layers of silicene.
Refer to caption(a)(b)
@@
Figure 7: (a) The change in energy as an atom is move across the vacuum from the bottom (110) surface of a silicon slab to the top (110) surface. (b) The z component of the force on the atom being moved is shown. Curves for different values of $\sigma$ are offset (black arrows) to aid comparison.